A sheet of cardboard of length 50 cm and width 26 cm has squares cut out of the corners so that it can be folded to form an open box with a base area of 640 cm2.
Find the dimensions of the cut out squares and sketch the obtained equation

Answer by edjones(8007)

Let x be one of the sides of a square.
(50-2x)(26-2x)=640
4x^2-152x+1300=640
x^2-38x+325=160
x^2-38x+165=0
(x-33)(x-5)=0
x=5cm
The dimensions:
base: 40 by 16
each long side: 5 by 40
Each short side: 5 by 16
.
Ed